OpenStudy (anonymous):

You are given an extended ratio that compares the lengths of the sides of a triangle. Find the length of each side. The ratio is 2:1:2. The longest side is 2x, the medium side is x+4, and the shortest side is 4.

OpenStudy (anonymous):

The given ratio: 2: 1: 2 The shortest side is 4. The medium is (x + 4) = 2 x 4 = 8 --> x = 8 - 4 = 4 The longest side is 2x = 2*4 = 8. The sides of the triangle are: 8: 4: 8 that satisfy the given ratio .